OpenStudy (anonymous):

Medal and fan reward!!! ^^ A baseball team has won 15 games and lost 9. If these games represent 16 2/3% of the games to be played, how many more games must the team win to average .750 for the season? a-63 b-75 c-93 d-100 e-108

OpenStudy (tkhunny):

(Total Wins)/(Total Games) = 0.750 15+9 = 24 24 is 16 2/3% of what?

OpenStudy (anonymous):

144...?

OpenStudy (tkhunny):

144/6 = 24 -- Correct. How many games must be won in total - forgetting that some already have been played?

OpenStudy (anonymous):

144?

OpenStudy (tkhunny):

No, only 75% of 144 is required. No one wins ALL their games!

OpenStudy (anonymous):

36?

OpenStudy (tkhunny):

What is 36? 36 is certainly not 75% of 144.

OpenStudy (anonymous):

Oh wait..I subtracted by accident. 108?

OpenStudy (tkhunny):

144*0.75 = 108 -- Correct. Now the question is, how many games must the win from the remaining games in order to get to 75%? They already have 15 wins and they need 108 total.

OpenStudy (anonymous):

93

OpenStudy (tkhunny):

How many games are left?

OpenStudy (tkhunny):

I ask this additional question only because there had better be at least 93 games remaining. If there are only 85 games remaining, it is VERY DIFFICULT to win 93 of them.

OpenStudy (anonymous):

Um...93..?

OpenStudy (tkhunny):

144-15-9 = 120 93/120 = 0.775 <== They need to win a lot of games from now on.